OpenStudy (anonymous):

solve the below equation for x. Round your answer to two decimal places. log(base 7) (1-x)=3

OpenStudy (raden):

convert to exponent form : 7^3 = 1 - x solve for x

OpenStudy (anonymous):

ok so it would be 343=1-x??

OpenStudy (raden):

yes, then solve for x

OpenStudy (anonymous):

I got -342????

OpenStudy (anonymous):

but that doesn't really make sense

OpenStudy (raden):

correct

OpenStudy (raden):

why not ?

OpenStudy (anonymous):

well I thought it was going to be a decimal or something cuz it said to round ur answer to 2 decimal places

OpenStudy (jdoe0001):

$$\bf \large { log_7 (1-x)=3\\ \text{using log cancellation rule of }\\ \color{blue}{a^{log_ax} =x}\\ 7^{log_7 (1-x)} = 7^3 \implies 1-x = 7^3 } $$

OpenStudy (raden):